# Quasi-Normal Modes of Loop Quantum Black Holes Formed from Gravitational Collapse

### v2

@@ II.2 Perturbations of a Scalar Field
Since in the rest of this paper we shall work only with the background described by the metric (2.10) and the dimensionless coordinates $\tilde{t}$ as well as $\tilde{r}$, without causing any confusion, we shall drop the tildes from the dimensionless coordinates $(\tilde{t},\tilde{r})$ and the corresponding quantities. Then, following Ramin2021 we consider the scalar perturbations that obey the Klein-Gordon equation
@@
Notice that, since $r$ is dimensionless, $x$ is actually also dimensionless. Also notice that, when the background has spherically symmetry, the perturbations are independent of the chociue of the angular index $m$ Chandra92 . Thus, without loss of the generality, we shall set $m=0$. For a given function $f(r)$ and choosing a suitable value of $l$, we can solve (2.12) for $\Psi$ (which contains the information of QNMs) for a BH.
@@ III Quasi-Normal Modes of LQBHs
First of all, we try to solve Eq.(2.12) in the time-domain with the finite difference method (FDM) Chao2023a ; Chao2022 ; RichardB . One advantage of such a method is that, for any chosen $l$, the final result from it can reflect comprehensively the influence of a bunch of different modes (represented by $\{m,n\}$, although $m$ can be always set to zero in the spherical case without loss of the generality, as mentioned earlier. It must not be confused this m with the mass parameter M of the LQBHs.). To apply the FDM, we first introduce two new variables $\mu\equiv t-x$ and $\nu\equiv t+x$ [so that $t=(\nu+\mu)/2$ and $x=(\nu-\mu)/2$]555 An important feature of the FDM is that in using it, we need the exact form of $x(r)$, in addition to its derivative with respect to $r$. Therefore, according to the definition (2.14), we have to assign $x$ an integral constant to absolutely fix it. In fact, such a constant could be chosen arbitrarily and it’s independent of our results for calculating QNMs. Thus, we made a simple choice by letting $x(r=2)=2$, which is consistent with the general choice in GR (See e.g., Leaver1985 ).. Therefore, on a $(N+1)\times(N+1)$ lattice (where $N$ is a positive integer that will be chosen properly according to our usage), we perform the calculation of $\Psi(\mu,\nu)$ by using the recursion formula

### v3

@@ II.2 Perturbations of a Scalar Field
Since in the rest of this paper we shall work only with the background described by the metric (2.10) and the dimensionless coordinates $\tilde{t}$ as well as $\tilde{r}$, without causing any confusion, we shall drop the tildes from the dimensionless coordinates $(\tilde{t},\tilde{r})$ and the corresponding quantities. Then, following Ramin2021 we consider the scalar perturbations that obey the Klein-Gordon equation  [^5] In this paper, we consider only the relativistic dispersion relation. In principle, higher-order corrections can be also included. But, these high-order terms are of order of ${\cal{O}}(K^{n}/M_{pl}^{n-2})$ with $n\geq 4$, where $K$ denotes the curvature of the macroscopic BHs, which is quite lower than the Planck scale, $M_{pl}\gg K$ outsides of such macroscopic BHs. Therefore, for the QNMs of macroscopic BHs, these corrections are negligible. For more details, see, for example, Ref. Wang2017 and references therein.
@@
Notice that, since $r$ is dimensionless, $x$ is actually also dimensionless. Also notice that, when the background has spherically symmetry, the perturbations are independent of the choice of the angular index $m$ Chandra92 . Thus, without loss of the generality, we shall set $m=0$. For a given function $f(r)$ and choosing a suitable value of $l$, we can solve (2.12) for $\Psi$ (which contains the information of QNMs) for a BH.
@@ III Quasi-Normal Modes of LQBHs
First of all, we try to solve Eq.(2.12) in the time-domain with the finite difference method (FDM) Chao2023a ; Chao2022 ; RichardB . One advantage of such a method is that, for any chosen $l$, the final result from it can reflect comprehensively the influence of a bunch of different modes (represented by $\{m,n\}$, although $m$ can be always set to zero in the spherical case without loss of the generality, as mentioned earlier. It must not be confused this m with the mass parameter M of the LQBHs.). To apply the FDM, we first introduce two new variables $\mu\equiv t-x$ and $\nu\equiv t+x$ [so that $t=(\nu+\mu)/2$ and $x=(\nu-\mu)/2$]666 An important feature of the FDM is that in using it, we need the exact form of $x(r)$, in addition to its derivative with respect to $r$. Therefore, according to the definition (2.14), we have to assign $x$ an integral constant to absolutely fix it. In fact, such a constant could be chosen arbitrarily and it’s independent of our results for calculating QNMs. Thus, we made a simple choice by letting $x(r=2)=2$, which is consistent with the general choice in GR (See e.g., Leaver1985 ).. Therefore, on a $(N+1)\times(N+1)$ lattice (where $N$ is a positive integer that will be chosen properly according to our usage), we perform the calculation of $\Psi(\mu,\nu)$ by using the recursion formula